C2H4 + O2 > CO2 + H2O

A. What is my theoretical yield of H2O if I start with 9.35 g of O2?

B. If I actually produced 2.63 g of H2O, what is my percent yield?
1
Expert's answer
2020-02-28T12:14:41-0500

9.35g Xg

C2H4 + 3O2 > 2CO2 + 2H2O  

3*32g 2*18g

A: m(theor H2O) = X = 9.35*2*18/(3*32) = 3.506g.

B: η(H2O) = m(theor H2O)/m(pract H2O) = 2.63g/3.506g*100% = 75.01% = 75%.